MATRIX ISOLATION AND COMPUTATIONAL STUDY OF iso-CF_{2}Br_{2}: A ROUTE TO MOLECULAR PRODUCTS IN CF_{2}Br_{2} PHOTOLYSIS

\underline{LISA GEORGE}, AIMABLE KALUME AND SCOTT A. REID, Department of Chemistry, Marquette University, Milwaukee, WI 53233; PATRICK Z. EL-KHOURY AND ALEXANDER TARNOVSKY, Department of Chemistry and Center for Photochemical Sciences, Bowling Green State University, Bowling Green, OH 43403.

The photolysis products of dibromodifluoromethane following selected wavelength laser irradiation were characterized by matrix isolation infrared and UV/Visible spectroscopy, supported by ab initio calculations. Photolysis at wavelengths of 240 and 266 nm of CF_{2}Br_{2}:Ar samples ( 1:5000) held at 5 K yielded iso-CF_{2}Br_{2} (F_{2}CBrBr), a weakly bound isomer of CF_{2}Br_{2}, which is characterized here for the first time. The observed infrared and UV/Visible absorptions of iso-CF_{2}Br_{2} are in excellent agreement with computational predictions at the B3LYP/aug-cc-pVTZ level. Single point energy calculations at the CCSD(T)/aug-cc-pVDZ level on the B3LYP optimized geometries show that the iso-form is a minimum on the CF_{2}Br_{2} potential energy surface, lying some 55 kcal/mol above the CF_{2}Br_{2} ground state. The energies of various stationary points on the CF_{2}Br_{2} PES were characterized computationally; taken with our experimental results, these show that iso-CF_{2}Br_{2} is an intermediate in the Br + CF_{2}Br reaction leading to molecular products (CF_{2} + Br_{2}). The photochemistry of the iso-form was also investigated; excitation into the intense 359 nm absorption band resulted in isomerization to CF_{2}Br_{2}. Our results are discussed in view of the rich literature on the gas-phase photochemistry of CF_{2}Br_{2}, particularly with respect to the existence of a roaming atom pathway leading to molecular products.